OpenStudy (anonymous):

Complete the equation based on the word problem then solve the equation for n. The temperature decreased 18°F between 10 p.m. and 4 a.m. The temperature at 4 a.m. was 69°F. What was the temperature at 10 p.m.?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@dmezzullo

OpenStudy (destinymasha):

n - 18 = 69? , so i would think you add.

OpenStudy (destinymasha):

69 + 18 = 87. so.... 87 - 18 degrees = 69 degrees